### v2

# Optimal alignment of Lorentz orientation and generalization to matrix Lie groups

Abstract
There exist elegant methods of aligning point clouds in $\mathbb{R}^{3}$. Unfortunately, these methods rely on the positive definite property of the Euclidean metric, and do not easily extend to the indefinite Minkowski metric. In this paper, we propose two solutions to the following problem: given inertial reference frames $A$ and $B$, and given (possibly noisy) measurements of a set of 4-vectors $\{v_{i}\}$ made in those reference frames with components $\{v_{A,i}\}$ and $\{v_{B,i}\}$, find the optimal Lorentz transformation $\Lambda$ such that $\Lambda v_{A,i}=v_{B,i}$. The method we outline is conceptually simple and easily extends to alignment problems in other matrix Lie groups.
reference frame alignment, Lorentz group
I Introduction
The problem of finding the optimal rigid alignment between two point clouds in $\mathbb{R}^{3}$ is well-known, and is also referred to as the orthogonal Procrustes problem. Two particularly elegant and efficient solutions exist, including the Kabsch algorithm [1] and the Horn algorithm [2]. These algorithms require as input two lists of vector components $\{v_{A,i}\}$ and $\{v_{B,i}\}$, which share a common origin, where $v_{A,i}$ is to be mapped to $v_{B,i}$. The Kabsch algorithm easily extends to arbitrary dimensions $N$, and produces an element of $SO(N)$. In contrast, the Horn algorithm relies on the geometric connection between unit quaternions and 3D rotations to rephrase the problem into a principal eigenvalue/eigenvector problem, and so it can only be used for $SO(3)$.
Both methods rely on the positive definiteness of the Euclidean metric, or equivalently that $SO(N)$ is compact, ensuring that the resulting optimization problem is convex. In contrast, the Minkowski inner product (we use the convention $\eta_{\mu\nu}=\text{diag}(-,+,+,+)$ in this work) induces a norm which is indefinite, and the proper orthochronous Lorentz group $SO(3,1)_{+}$ is noncompact due to the Lorentz boosts. Our goal is to find an efficient algorithm for performing alignment.
An alternative approach is needed to find the optimal Lorentz transformation from one set of 4-vectors to another. In this work, we present two methods of performing this alignment. The first method is traditional least squares, constrained optimization using Newton-Raphson iteration, which has the advantage of being highly numerically robust, but is computationally intensive. The second method is by solving for an least squares linear transformation with the Moore-Penrose pseudoinverse (which does not guarantee a Lorentz matrix such that $L\eta L^{T}=\eta$), and projecting that linear transformation onto the nearest element of the Lorentz group via a detour through the Lie algebra $\mathfrak{so}(3,1)$.
II Insights into pose alignment from the Kabsch and Horn algorithms
Note that in flat space, points and tangent vectors are conflated. In this work, we assume that all inputs are true tangent vectors, and there is no “centering” process required.
II.1 The Kabsch algorithm
The proof proceeds as follows: for two lists $\{v_{A,i}\}$ and $\{v_{B,i}\}$ of $n$ vectors in $N$ dimensions, let $A$ and $B$ be $N\times n$ matrices where $A_{ji}=(v_{A,i})_{j}$ and $B_{ji}=(v_{B,i})_{j}$. We seek a rotation matrix $R$ such that $Rv_{A,i}\approx v_{B,i}$.
First, we write the objective function:
$\sum_{i}\|v_{B,i}-Rv_{A,i}\|^{2}=\|B-RA\|^{2}$ (1)
$=\text{tr}\left((B-RA)^{T}(B-RA)\right)$ (2)
$=\text{tr}(B^{T}B)-2\text{tr}(A^{T}R^{T}B)+\text{tr}(A^{T}A).$ (3)
In Eq (1), the left side norm is the least squares (i.e. Frobenius) norm for vectors, and the norm on the right side is the least squares norm for matrices. In Eq (3), the first and final terms are independent of $R$, we may drop them and seek to minimize the middle term, or after using the cyclicity property of the trace, maximize:
$f(R)=\text{tr}(R^{T}BA^{T}).$ (4)
The key insight in the Kabsch algorithm is to perform singular value decomposition on the matrix $H=BA^{T}=U\Sigma V^{T}$, where $U,V\in O(N)$, and $\Sigma=\text{diag}(\sigma_{1},\sigma_{2},\cdots)$ contains the singular values.
$f(R)=\text{tr}(R^{T}H)$ (5)
$=\text{tr}(R^{T}U\Sigma V^{T})$ (6)
$=\text{tr}(V^{T}R^{T}U\Sigma).$ (7)
Since $V^{T},R^{T},U\in O(N)$, the resulting product is $M=V^{T}R^{T}U\in O(N)$. Furthermore, $|M_{ii}|\leq 1$ and $\text{tr}(M_{ii})\leq N$.
$Tr(M\Sigma)=\sum_{i}\sigma_{i}M_{ii}$ (8)
This expression is maximized under the conditions above when $M$ is the identity.
$M=V^{T}R^{T}U=I$ (9)
$\implies R=UV^{T}$ (10)
To ensure that $R$ is a proper rotation, we require $|R|=1$. When $|UV^{T}|=-1$, we may negate the last column of $U$ to achieve this condition.
II.2 Quaternions
Quaternions can be thought of as a four-dimensional real vector space with basis $\{\mathbf{1},\mathbf{i},\mathbf{j},\mathbf{k}\}$, with quaternion addition defined by componentwise addition. We will represent quaternions as $q=(q_{0},q_{1},q_{2},q_{3})$, where $q_{0},q_{1},q_{2},q_{3}\in\mathbb{R}$. We denote the imaginary unit in $\mathbb{C}$ by $h^{2}=-1$. The operation of quaternion conjugation is defined by
$q^{*}=(q_{0},-q_{1},-q_{2},-q_{3}),$ (11)
Further algebraic properties define the operation of quaternion multiplication:
$\mathbf{i}\mathbf{j}=\mathbf{k},$ (12)
$\mathbf{j}\mathbf{k}=\mathbf{i},$ (13)
$\mathbf{k}\mathbf{i}=\mathbf{j},$ (14)
$\mathbf{i}^{2}=\mathbf{j}^{2}=\mathbf{k}^{2}=-\mathbf{1},$ (15)
so that
$(p_{0},p_{1},p_{2},p_{3})$ $(q_{0},q_{1},q_{2},q_{3})$
=(=(= ( $p_{0}q_{0}-p_{1}q_{1}-p_{2}q_{2}-p_{3}q_{3},$
$p_{0}q_{1}+p_{1}q_{0}+p_{2}q_{3}-p_{3}q_{2},$
$p_{0}q_{2}-p_{1}q_{3}+p_{2}q_{0}+p_{3}q_{1},$
$p_{0}q_{3}+p_{1}q_{2}-p_{2}q_{1}+p_{3}q_{0}).$ (16)
Considering $q$ as a column vector, quaternion multiplication $pq$ can also be described as left multiplication of that column vector by a matrix
$pq=\begin{bmatrix}p_{0}&-p_{1}&-p_{2}&-p_{3}\\ p_{1}&p_{0}&p_{3}&-p_{2}\\ p_{2}&-p_{3}&p_{0}&p_{1}\\ p_{3}&p_{2}&-p_{1}&p_{0}\end{bmatrix}\begin{bmatrix}q_{0}\\ q_{1}\\ q_{2}\\ q_{3}\end{bmatrix}.$ (17)
This interpretation of quaternion multiplication as matrix multiplication is used extensively by Horn. Quaternion multiplication is associative but not commutative.
The quaternion norm is defined as $||q||_{\text{quat}}=qq^{*}=q_{0}^{2}+q_{1}^{2}+q_{2}^{2}+q_{3}^{2}$, and unit quaternions by $||q||_{\text{quat}}=1$. It is unit quaternions which are isomorphic $SU(2)$, which is a double cover of $SO(3)$. To represent rotations, a vector $\mathbf{r}\in\mathbb{R}^{3}$ is encoded as a quaternion $r=(0,r_{1},r_{2},r_{3})$, and rotation by a matrix $R\in SO(3)$ is accomplished in quaternion space by:
$R\mathbf{r}\sim qrq^{*},$ (18)
where $q$ is the quaternion representing $R$, the product $R\mathbf{R}$ is a matrix multiplication, and $qrq^{*}$ is two quaternion multiplications.
II.3 Horn’s algorithm
We will not repeat the entirety of Horn’s derivation here [2]. We adapt our notation here to match the preceding sections, so that $\mathbf{v}_{A,i}$ and $\mathbf{v}_{B,i}$ are ordinary vectors in $\mathbb{R}^{3}$, and $v_{A,i}$ and $v_{B,i}$ are their quaternion counterparts using Eq (18). Instead of minimizing the difference between vector components, Horn tries to maximize the dot products of the rotated first set of vectors with the second set of vectors, i.e. the equivalent of Eq (4) from the proof of Kabsch’s algorithm.
$f(q)=\sum_{i}(qv_{A,i}q^{*})\cdot v_{B,i}.$ (19)
In this expression, $q$ is a unit quaternion, and $\cdot$ is the true Euclidean inner product. Using the interpretation of quaternion multiplication as a matrix multiplication, Horn switches notation once again, rewriting $f(q)$ so that it becomes a quadratic form:
$f(q)=\mathbf{q}^{T}N\mathbf{q}.$ (20)
In this expression, $\mathbf{q}$ are the components of the quaternion $q$, now regarded as a column vector. $N$ is a symmetric matrix formed using the entries in $H=BA^{T}$ from the Kabsch derivation. The key insight now is that Eq (20) can be analyzed in terms of the eigendecomposition of $N$. N is symmetric, and therefore it possesses a complete, orthonormal eigenbasis. Since the eigenvectors ${e_{i}}$ are a basis for $\mathbb{R}^{4}$ the desired unit quaternion may be expressed as $\mathbf{q}=a_{1}\mathbf{e}_{1}+a_{2}\mathbf{e}_{2}+a_{3}\mathbf{e}_{3}+a_{4}\mathbf{e}_{4}$, with $a_{i}\in\mathbb{R}$. The norm $||q||_{\text{quat}}$ can be calculated as $a_{1}^{2}+a_{2}^{2}+a_{3}^{2}+a_{4}^{2}=1$, and now:
$\mathbf{q}^{T}N\mathbf{q}=\alpha_{1}^{2}\lambda_{1}+\alpha_{2}^{2}\lambda_{2}+\alpha_{3}^{2}\lambda_{3}+\alpha_{4}^{2}\lambda_{4},$ (21)
which is maximized when the largest eigenvalue $\lambda_{\text{max}}$ and its corresponding eigenvector $\mathbf{e}_{\text{max}}$ are chosen as $\mathbf{q}$. One can therefore use a power iteration method or similar algorithm to calculate the principal eigenvalue and eigenvector of $N$ [3].
II.4 Can we use the Kabsch or Horn algorithms as inspiration for Minkowski vector alignment?
If we consider minimizing the Frobenius norm of the components, then we follow the Kabsch algorithm in defining the objective function as:
$f(\Lambda)=\sum_{i}\|v_{B,i}-\Lambda v_{A,i}\|^{2}.$ (22)
In the Kabsch algorithm, singular value decomposition automatically produces elements of $O(N)$, which allows the simplification step in Eq (7) to Eq (8). Further progress using the Kabsch approach would require an efficient numerical Lorentz singular value decomposition [4], which to our knowledge does not exist.
In the Horn algorithm, the situation becomes more complex (literally). Just as unit quaternions represent rotations in $SO(3)$, unit biquaternions represent Lorentz transformations [5]. A biquaternion is a vector space over $\mathbb{C}$ rather than $\mathbb{R}$. The rules of quaternion conjugation and multiplication remain identical. Crucially, whereas vectors in $\mathbb{R}^{3}$ were packaged into the imaginary part of the quaternion, 4-vectors must instead be packaged as Minkowski biquaternions, also known as minquats, so that a 4-vector $\mathbf{v}=(t,x,y,z)$ is represented as
$v=(t,ix,iy,iz).$ (23)
Biquaternions have two types of conjugation: complex conjugation in the field, defined by
$\bar{q}=(\bar{q}_{0},\bar{q}_{1},\bar{q}_{2},\bar{q}_{3}),$ (24)
and conjugation over the quaternion algebra, represented by $q^{*}$ in Eq (11). The two types of conjugation commute $(\bar{q})^{*}=\overline{(q^{*})}=\overline{q^{*}}$. The biquaternion norm is defined as the complex number $||q||_{\text{biquat}}=qq^{*}\in\mathbb{C}$. The action of the unit quaternion $q$ associated with Lorentz transformation $\Lambda$ on a minquat $v$ is defined by
$\Lambda_{q}(v)=qv\overline{q^{*}}.$ (25)
One may postulate that perhaps we can maximize a Minkowski inner product (for alignment of time-like vectors), since many of the quaternion properties still hold. Unfortunately, a unit biquaternion’s components may be unbounded, for example $(i\cosh\phi,\sinh\phi,0,0)$ which has biquaternion norm $\sinh^{2}\phi-\cosh^{2}\phi=1$. The presence of negative contributions to the norm prevents us from writing down the maximum as in Eq (21).
Schematically, instead of maximizing a quadratic form, we would be maximizing a Hermitian form $\overline{\mathbf{q}}^{T}N\mathbf{q}$, where $N$ is Hermitian and therefore has real eigenvalues $\lambda_{i}$ and orthogonal eigenvectors $e_{i}$, and so that $q$ must be a complex linear combination of the $e_{i}$:
$q=\alpha_{1}e_{1}+\alpha_{2}e_{2}+\alpha_{3}e_{3}+\alpha_{4}e_{4},$ (26)
and we must maximize
$\overline{\mathbf{q}}^{T}N\mathbf{q}=|\alpha_{1}|^{2}\lambda_{1}+|\alpha_{2}|^{2}\lambda_{2}+|\alpha_{3}|^{2}\lambda_{3}+|\alpha_{4}|^{2}\lambda_{4}$ (27)
subject to the condition
$\alpha_{1}^{2}+\alpha_{2}^{2}+\alpha_{3}^{2}+\alpha_{4}^{2}=1$ (28)
We emphasize that there is no typo in omitting the absolute value symbols in Eq (28), since the biquaternion norm can be complex. Unlike in the Horn derivation, we have not linearized the problem at all, and have merely turned one constrained nonlinear optimization problem into another.
III Proposed solutions to the Lorentz alignment problem
III.1 Review of the Lorentz algebra and group
As a reminder, matrix Lie groups such as $SO(3,1)_{+}$ are related to a Lie algebra, e.g. $\mathfrak{so}(3,1)$, via the matrix logarithm and matrix exponential functions. The Lorentz algebra is the set of matrices:
$A=\begin{bmatrix}0&\zeta^{1}&\zeta^{2}&\zeta^{3}\\ \zeta^{1}&0&-\theta^{3}&\theta^{2}\\ \zeta^{2}&\theta^{3}&0&-\theta^{1}\\ \zeta^{3}&-\theta^{2}&\theta^{1}&0\end{bmatrix}\in\mathfrak{so}(3,1),$ (29)
where $\bm{\zeta}=[\zeta^{1},\zeta^{2},\zeta^{3}]$ is the boost vector and $\bm{\theta}=[\theta^{1},\theta^{2},\theta^{3}]$ is the rotation vector, and both are real. The matrix exponential of this set of matrices (i.e. the Lorentz analogue of Rodrigues’ rotation formula [6]) has recently been explicitly computed [7]:
$\Lambda(\bm{\zeta},\bm{\theta})=\exp A=\frac{1}{a^{2}+b^{2}}\sum_{i=0}^{3}f_{i}(a,b)A^{i},$ (30)
$a^{2}=\frac{1}{2}\left[|\bm{\theta}|^{2}-|\bm{\zeta}|^{2}+\sqrt{\left(|\bm{\theta}|^{2}-|\bm{\zeta}|^{2}\right)^{2}+4\left(\bm{\theta}\cdot\bm{\zeta}\right)^{2}}\right],$ (31)
$b^{2}=\frac{1}{2}\left[|\bm{\zeta}|^{2}-|\bm{\theta}|^{2}+\sqrt{\left(|\bm{\theta}|^{2}-|\bm{\zeta}|^{2}\right)^{2}+4\left(\bm{\theta}\cdot\bm{\zeta}\right)^{2}}\right],$ (32)
$f_{0}(a,b)=b^{2}\cos a+a^{2}\cosh b,$ (33)
$f_{1}(a,b)=\frac{b^{2}}{a}\sin a+\frac{a^{2}}{b}\sinh b,$ (34)
$f_{2}(a,b)=\cosh b-\cos a,$ (35)
$f_{3}(a,b)=\frac{\sinh b}{a}-\frac{\sin a}{a}.$ (36)
Haber notes that Eqs (33)-(35) are independent of the overall signs of $a$ and $b$, which is evident by noting that $\cos$ and $\cosh$ are even, whereas $\sin$ and $\sinh$ are odd. Since it is a continuous function, the matrix exponential reproduces only the proper orthochronous part of $SO(3,1)$, which is what we would presumably desire of an alignment algorithm.
III.2 Method 1: Direct nonlinear optimization
For naive least squares optimization of alignment, we write down the objective function:
$f(\bm{\zeta},\bm{\theta})=\sum_{i}\|v_{B,i}-\Lambda(\bm{\zeta},\bm{\theta})v_{A,i}\|^{2},$ (37)
which can be solved either through gradient descent or other method [8]. We opt to use SciPy’s general solvers for this task [9]. We will refer to this method as the direct minimization method.
III.3 Method 2: Projection onto the Lorentz group after unconstrained solving
A common method to solve problems of the form
$Ax=b$ (38)
in a least squares sense is to use the Moore-Penrose inverse (or pseudoinverse) $A^{+}$:
$x=A^{+}b,$ (39)
where $A^{+}=(A^{T}A)^{-1}A^{T}$ when $(A^{T}A)^{-1}$ exists [8]. Consider now the $4\times n$ matrices $X_{ij}=(v_{A,j})_{i}$ and $Y_{ij}=(v_{B,j})_{i}$ (rows index the dimension, and columns index the vectors to be aligned), and the equation
$\Lambda_{0}X=Y.$ (40)
The Moore-Penrose inverse gives a least squares estimate for a linear transformation $\Lambda_{0}$ which transforms the vectors $v_{A,j}$ to $v_{B,j}$, but which is not necessarily a Lorentz transform. In other words, we assume that there is some measurement error in $X$ and $Y$ which prevents $\Lambda_{0}$ from satisfying $\Lambda_{0}^{T}\eta\Lambda_{0}=\eta$. To project $\Lambda_{0}$ onto a nearby true Lorentz transformation, we take a matrix logarithm:
$l_{0}=\log\Lambda_{0}=\begin{bmatrix}m_{11}&m_{12}&m_{13}&m_{14}\\ m_{21}&m_{22}&m_{23}&m_{24}\\ m_{31}&m_{32}&m_{33}&m_{34}\\ m_{41}&m_{42}&m_{43}&m_{44}\end{bmatrix},$ (41)
and force $l_{0}$ to be of the form given in Eq (29). Specifically, (1) we set the diagonal of $l_{0}$ to be 00, (2) symmetrize the first row and first column, and (3) antisymmetrize the bottom right $3\times 3$ block:
$l=\frac{1}{2}\begin{bmatrix}0&m_{12}+m_{21}&m_{13}+m_{31}&m_{14}+m_{41}\\ m_{12}+m_{21}&0&m_{23}-m_{32}&m_{24}-m_{42}\\ m_{13}+m_{31}&m_{32}-m_{23}&0&m_{34}-m_{43}\\ m_{14}+m_{41}&m_{42}-m_{24}&m_{43}-m_{34}&0\end{bmatrix}$ (42)
Finally, we take $\Lambda=\exp l$ using Eqs (30)-(35) as the optimal Lorentz transformation. The objective function which this procedure minimizes is:
$f(l)=\|l-l_{0}\|^{2},$ (43)
$l\in\mathfrak{so}(3,1)$ (44)
or
$f(\Lambda)=\|\log\left(\Lambda^{-1}\Lambda_{0}\right)\|^{2},$ (45)
$\Lambda\in SO(3,1).$ (46)
This method of projection extends easily to all other matrix Lie groups: one simply proposes and minimizes an objective function on the Lie algebra. Extra consideration is needed if the Lie group is not simply connected and one desires elements other than from the connected component containing the identity. We refer to this method as the Lie algebra method.
IV Results
We implemented both methods in Python, benchmarking accuracy and timing. We used NumPy [10] for array operations, and SciPy [9] matrix logarithms, matrix exponentials, and nonlinear solving. All testing was performed on an M3 MacBook Air.
IV.1 Ideal conditions with no noise
We began with a set of four time-like 4-vectors:
$X=\begin{bmatrix}1&0&0&0\\ \sqrt{2}&1&0&0\\ \sqrt{2}&0&1&0\\ \sqrt{2}&0&0&1\end{bmatrix}$ (47)
and a simple boost along $x$ of the form
$\Lambda=\begin{bmatrix}\gamma&-\beta\gamma&0&0\\ -\beta\gamma&\gamma&0&0\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix},$ (48)
with $\beta=0.3$ and $\gamma=\frac{1}{\sqrt{1-\beta^{2}}}$, and verified that both methods could reproduce $\Lambda$, given $A$ and $B=A\Lambda^{T}$. This simple sanity check showed that both methods indeed work. The use of the explicit form of the Lorentz boost given by Haber in Eqs (30)-(35) ensured that the resulting matrix indeed had determinant close to $1$. However, the Lie algebra method was slightly more effective than direct least squares optimization via the default solver in scipy.optimize.minimize in ensuring that the components of $A\Lambda_{\text{estimated}}^{T}$ were close to $A\Lambda^{T}$, as demonstrated in the Supplemental Materials. For the Lie algebra method, we tested brute force matrix exponentiation via scipy.linalg.expm compared to Haber’s explicit computation, and found that both computations resulted in $\det\Lambda\approx 1$ to 14 decimal places.
The timing between the two Lie algebra methods was also extremely similar, with each iteration taking 700-800 $\mu$s, with standard deviation of 20-30 $\mu$s. In contrast, the default solver for scipy.optimize.minimize took an average of 24 ms, which was 30$\sim$ 30$\times$ 30 × as slow as the Lie algebra method.
IV.2 Varying the number of vectors given, noise, and averaging over arbitrary Lorentz transformations
Figure 1: The Frobenius ($L^{2}$) norm of the estimated Lorentz matrix from frame $A$ to frame $B$ as compared to the actual Lorentz matrix. The rows represent varying the noise $\epsilon$ present in the measurement of the vectors in frame $B$. The columns represent varying the number $n$ of vectors upon which the estimation was performed. Evidently, the Lie algebra method and the direct minimization method are equivalent in accuracy across a range of testing conditions.
Note that the joint distribution for $\bm{\zeta}$ and $\bm{\theta}$ is a spherically symmetric product of normal distributions. We generated $n$ random 4-vectors with Minkowski norm $-1$ by sampling $x,y,z\sim\mathcal{N}(0,0.3)$ and setting $t=\sqrt{1+x^{2}+y^{2}+z^{2}}$, and used these vectors as our $v_{A,i}$. We introduced noise in the vector components as measured in frame $B$ by adding random noise $\Delta x,\Delta y,\Delta z\sim\mathcal{N}(0,\epsilon)$ to $x,y,z$, normalizing the time component for $x+\Delta x,y+\Delta y,z+\Delta z$ again, and then boosting by $\Lambda$ to these vectors.
We observe that the overall distribution of both the maximum element-wise error and the overall $L^{2}$ error is quite similar between the two proposed methods (Figs 1, 2).
Figure 2: The max ($L^{\infty}$) norm of the estimated Lorentz matrix from frame $A$ to frame $B$ as compared to the actual Lorentz matrix. The rows and columns represent the same variables as in Fig 1.
In Figs 1and 2, we compare the Frobenius and max norms of the estimated Lorentz transformation matrix against the true Lorentz transformation. We generated $1,000$ random Lorentz transformations using
$\zeta^{i}\sim\mathcal{N}(0,0.2),$ (49)
$\theta^{i}\sim\mathcal{N}(0,1).$ (50)
V Discussion
In this work, we have proposed two methods of solving the Lorentz alignment problem. First, we showed that the principles behind the Kabsch algorithm and Horn algorithm for rigid rotation alignment in Euclidean $\mathbb{R}^{3}$ do not easily extend to the Lorentz geometry. A Kabsch-like approach is not feasible unless a Lorentz singular value decomposition is available, the development of which is beyond the scope of this work. The Horn-like approach is infeasible because of the unbounded nature of the unit biquaternions representation of the Lorentz group, as compared to the bounded real quaternion representation of $SO(3)$; this obstacle prevents the optimization problem from linearizing into a principal eigenvalue/eigenvector problem.
Instead, we can either perform constrained optimization using gradient descent or related methods (the minimization method) or project the least-squares linear operator via the pseudoinverse onto the Lorentz group via a second least squares minimization in the Lie algebra. We have shown that at least in Python, the second approach is more performant, and better reproduces the answer in the case that an exact solution exists. We did not vary the minimization method used by SciPy’s default solver, though ideally an optimal solution would not be dependent on such as choice. The explicit formula relating the boost and rotation vectors to the corresponding Lorentz transformation was necessary in the minimization method, whereas it was optional in the Lie algebra method with little performance difference between it and brute force matrix exponentiation.
The Lie algebra method presented here technically also solves the orthogonal Procrustes problem by restricting our attention to the bottom right $3\times 3$ matrix. However, it extends to arbitrary matrix Lie groups, unlike the Kabsch algorithm which only works for $O(N)$, and the Horn algorithm which relies on a quaternion representation with bounded components. Therefore we feel that this work is valuable in the overall scheme of manifold alignment. An application of Lorentz alignment is finding the connection between free-falling reference frames in smooth lattice general relativity [11].
Future directions on this topic may provide more extensive benchmarking across other matrix Lie groups, and implement the methods described here in a lower level language.
Code and data availability
All code needed to reproduce the data and figures in this work are included in the Supplemental Materials as a Jupyter notebook.
Acknowledgments
The author would like to thank the Penn State College of Medicine’s Medical Scientist Training Program for their support of his career.
Conflict of interest statement
The author reports no potential conflicts of interest.

### v3

@@
 There exist elegant methods of aligning point clouds in $\mathbb{R}^{3}$.
-Unfortunately, these methods rely on the positive definite property of the Euclidean metric, and do not easily extend to the indefinite Minkowski metric.
-In this paper, we propose two solutions to the following problem: given inertial reference frames $A$ and $B$, and given (possibly noisy) measurements of a set of 4-vectors $\{v_{i}\}$ made in those reference frames with components $\{v_{A,i}\}$ and $\{v_{B,i}\}$, find the optimal Lorentz transformation $\Lambda$ such that $\Lambda v_{A,i}=v_{B,i}$.
-The method we outline is conceptually simple and easily extends to alignment problems in other matrix Lie groups.
+Unfortunately, these methods fail to generalize to the case of Minkowski space, as we will show.
+Instead, we propose two solutions to the following problem: given inertial reference frames $A$ and $B$, and given (possibly noisy) measurements of a set of 4-vectors $\{v_{i}\}$ made in those reference frames with components $\{v_{A,i}\}$ and $\{v_{B,i}\}$, find the optimal Lorentz transformation $\Lambda$ such that $\Lambda v_{A,i}=v_{B,i}$.
+The first method is direct least squares optimization through a parametrization of $SO(3,1)_{+}$ in terms of the familiar boost and rotation vectors.
+The second method takes a detour through the Lorentz algebra; in addition to being conceptually simple and possessing a computational advantage over the first method, it can easily be generalized to the alignment of vector representations in other matrix Lie groups.
 reference frame alignment, Lorentz group
@@
 The problem of finding the optimal rigid alignment between two point clouds in $\mathbb{R}^{3}$ is well-known, and is also referred to as the orthogonal Procrustes problem.
-Two particularly elegant and efficient solutions exist, including the Kabsch algorithm [1] and the Horn algorithm [2].
+Two particularly elegant and efficient solutions exist: the Kabsch algorithm [1] and the Horn algorithm [2].
 These algorithms require as input two lists of vector components $\{v_{A,i}\}$ and $\{v_{B,i}\}$, which share a common origin, where $v_{A,i}$ is to be mapped to $v_{B,i}$.
@@
 In contrast, the Minkowski inner product (we use the convention $\eta_{\mu\nu}=\text{diag}(-,+,+,+)$ in this work) induces a norm which is indefinite, and the proper orthochronous Lorentz group $SO(3,1)_{+}$ is noncompact due to the Lorentz boosts.
-Our goal is to find an efficient algorithm for performing alignment.
-An alternative approach is needed to find the optimal Lorentz transformation from one set of 4-vectors to another.
+First, we show that neither of the algorithms above can be applied directly to the problem of alignment in $SO(3,1)_{+}$, and an alternative approach is needed to find the optimal Lorentz transformation from one set of 4-vectors to another.
 In this work, we present two methods of performing this alignment.
@@
 The second method is by solving for an least squares linear transformation with the Moore-Penrose pseudoinverse (which does not guarantee a Lorentz matrix such that $L\eta L^{T}=\eta$), and projecting that linear transformation onto the nearest element of the Lorentz group via a detour through the Lie algebra $\mathfrak{so}(3,1)$.
+To our knowledge, this problem has not been addressed previously in the literature.
 II Insights into pose alignment from the Kabsch and Horn algorithms
-Note that in flat space, points and tangent vectors are conflated. In this work, we assume that all inputs are true tangent vectors, and there is no “centering” process required.
-II.1 The Kabsch algorithm
-The proof proceeds as follows: for two lists $\{v_{A,i}\}$ and $\{v_{B,i}\}$ of $n$ vectors in $N$ dimensions, let $A$ and $B$ be $N\times n$ matrices where $A_{ji}=(v_{A,i})_{j}$ and $B_{ji}=(v_{B,i})_{j}$.
-We seek a rotation matrix $R$ such that $Rv_{A,i}\approx v_{B,i}$.
-First, we write the objective function:
-$\sum_{i}\|v_{B,i}-Rv_{A,i}\|^{2}=\|B-RA\|^{2}$ (1)
-$=\text{tr}\left((B-RA)^{T}(B-RA)\right)$ (2)
+In this section, we first review the proofs of the Kabsch and Horn algorithms, and justify why neither method is directly applicable to alignment of Lorentz group.
+Note that in flat space, points and tangent vectors are conflated.
+In this work, we assume that all inputs are true tangent vectors, and there is no “centering” process required.
+II.1 The Kabsch and Horn algorithms
+Theorem 1.
+(Kabsch algorithm) Let there be two lists $\{\mathbf{v}_{A,i}\}$ and $\{\mathbf{v}_{B,i}\}$ of $n$ vectors in $N$ dimensions.
+Let $A$ and $B$ be $N\times n$ matrices where $A_{ji}=(\mathbf{v}_{A,i})_{j}$ and $B_{ji}=(\mathbf{v}_{B,i})_{j}$.
+Let the objective function be
+$L(R)=\sum_{i}\|\mathbf{v}_{B,i}-R\mathbf{v}_{A,i}\|^{2}=\|B-RA\|^{2},$ (1)
+where $\|\cdot\|$ is the Frobenius norm. Given the singular value decomposition of $BA^{T}=U\Sigma V^{T}$. If $|UV^{T}|=1$, the rotation matrix $R$ which minimizes $L(R)$ is
+$R=UV^{T},$ (2)
+and otherwise
+$R=U\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&-1\end{bmatrix}V^{T}.$ (3)
+Proof.
+First, we simplify the objective function:
+$L(R)=\text{tr}\left((B-RA)^{T}(B-RA)\right)$ (4)
 $=\text{tr}(B^{T}B)-2\text{tr}(A^{T}R^{T}B)+\text{tr}(A^{T}A).$ (5)
@@
 Frobenius) norm for vectors, and the norm on the right side is the least squares norm for matrices.
-In Eq (3), the first and final terms are independent of $R$, we may drop them and seek to minimize the middle term, or after using the cyclicity property of the trace, maximize:
+In Eq (5), the first and final terms are independent of $R$, we may drop them and seek to minimize the middle term, or equivalently after using the cyclicity property of the trace, maximize:
 $f(R)=\text{tr}(R^{T}BA^{T}).$ (6)
@@
 Since $V^{T},R^{T},U\in O(N)$, the resulting product is $M=V^{T}R^{T}U\in O(N)$. Furthermore, $|M_{ii}|\leq 1$ and $\text{tr}(M_{ii})\leq N$.
-$Tr(M\Sigma)=\sum_{i}\sigma_{i}M_{ii}$ (8)
+$\text{tr}(M\Sigma)=\sum_{i}\sigma_{i}M_{ii}$ (10)
 This expression is maximized under the conditions above when $M$ is the identity.
@@
 $\implies R=UV^{T}$ (12)
-To ensure that $R$ is a proper rotation, we require $|R|=1$. When $|UV^{T}|=-1$, we may negate the last column of $U$ to achieve this condition.
-II.2 Quaternions
-Quaternions can be thought of as a four-dimensional real vector space with basis $\{\mathbf{1},\mathbf{i},\mathbf{j},\mathbf{k}\}$, with quaternion addition defined by componentwise addition.
-We will represent quaternions as $q=(q_{0},q_{1},q_{2},q_{3})$, where $q_{0},q_{1},q_{2},q_{3}\in\mathbb{R}$.
-We denote the imaginary unit in $\mathbb{C}$ by $h^{2}=-1$.
-The operation of quaternion conjugation is defined by
-$q^{*}=(q_{0},-q_{1},-q_{2},-q_{3}),$ (11)
-Further algebraic properties define the operation of quaternion multiplication:
-$\mathbf{i}\mathbf{j}=\mathbf{k},$ (12)
-$\mathbf{j}\mathbf{k}=\mathbf{i},$ (13)
-$\mathbf{k}\mathbf{i}=\mathbf{j},$ (14)
-$\mathbf{i}^{2}=\mathbf{j}^{2}=\mathbf{k}^{2}=-\mathbf{1},$ (15)
-so that
-$(p_{0},p_{1},p_{2},p_{3})$ $(q_{0},q_{1},q_{2},q_{3})$
-=(=(= ( $p_{0}q_{0}-p_{1}q_{1}-p_{2}q_{2}-p_{3}q_{3},$
-$p_{0}q_{1}+p_{1}q_{0}+p_{2}q_{3}-p_{3}q_{2},$
-$p_{0}q_{2}-p_{1}q_{3}+p_{2}q_{0}+p_{3}q_{1},$
-$p_{0}q_{3}+p_{1}q_{2}-p_{2}q_{1}+p_{3}q_{0}).$ (16)
-Considering $q$ as a column vector, quaternion multiplication $pq$ can also be described as left multiplication of that column vector by a matrix
-$pq=\begin{bmatrix}p_{0}&-p_{1}&-p_{2}&-p_{3}\\ p_{1}&p_{0}&p_{3}&-p_{2}\\ p_{2}&-p_{3}&p_{0}&p_{1}\\ p_{3}&p_{2}&-p_{1}&p_{0}\end{bmatrix}\begin{bmatrix}q_{0}\\ q_{1}\\ q_{2}\\ q_{3}\end{bmatrix}.$ (17)
-This interpretation of quaternion multiplication as matrix multiplication is used extensively by Horn. Quaternion multiplication is associative but not commutative.
-The quaternion norm is defined as $||q||_{\text{quat}}=qq^{*}=q_{0}^{2}+q_{1}^{2}+q_{2}^{2}+q_{3}^{2}$, and unit quaternions by $||q||_{\text{quat}}=1$.
-It is unit quaternions which are isomorphic $SU(2)$, which is a double cover of $SO(3)$.
-To represent rotations, a vector $\mathbf{r}\in\mathbb{R}^{3}$ is encoded as a quaternion $r=(0,r_{1},r_{2},r_{3})$, and rotation by a matrix $R\in SO(3)$ is accomplished in quaternion space by:
-$R\mathbf{r}\sim qrq^{*},$ (18)
-where $q$ is the quaternion representing $R$, the product $R\mathbf{R}$ is a matrix multiplication, and $qrq^{*}$ is two quaternion multiplications.
-II.3 Horn’s algorithm
+To ensure that $R$ is a proper rotation, we require $|R|=1$. When $|UV^{T}|=-1$, we may negate the last column of $U$ to achieve this condition. ∎
+Theorem 2.
+(Horn algorithm) Given the same inputs as in Theorem 1, the unit quaternion corresponding to the rotation which minimizes $L(R)$ is the maximal vector of the symmetric matrix $N$, defined in Eq (14).
+Proof.
 We will not repeat the entirety of Horn’s derivation here [2].
+For a review of the quaternion algebra, refer to §A.
 We adapt our notation here to match the preceding sections, so that $\mathbf{v}_{A,i}$ and $\mathbf{v}_{B,i}$ are ordinary vectors in $\mathbb{R}^{3}$, and $v_{A,i}$ and $v_{B,i}$ are their quaternion counterparts using Eq (61).
-Instead of minimizing the difference between vector components, Horn tries to maximize the dot products of the rotated first set of vectors with the second set of vectors, i.e.
-the equivalent of Eq (4) from the proof of Kabsch’s algorithm.
+Horn attempts to maximize the dot products of the rotated first set of vectors with the second set of vectors, i.e. the equivalent of Eq (6) from the proof of the Kabsch algorithm.
 $f(q)=\sum_{i}(qv_{A,i}q^{*})\cdot v_{B,i}.$ (13)
-In this expression, $q$ is a unit quaternion, and $\cdot$ is the true Euclidean inner product.
-Using the interpretation of quaternion multiplication as a matrix multiplication, Horn switches notation once again, rewriting $f(q)$ so that it becomes a quadratic form:
+In this expression, $q$ is a unit quaternion, $qv_{A,i}q^{*}$ is computed via the quaternion product (Eq (59), and $\cdot$ is the Euclidean inner product.
+Using the interpretation of quaternion multiplication as a matrix multiplication (as in Eq (60)), Horn switches notation, rewriting $f(q)$ so that it becomes a quadratic form:
 $f(q)=\mathbf{q}^{T}N\mathbf{q}.$ (14)
@@
 Since the eigenvectors $e_{i}$} are a basis for $\mathbb{R}^{4}$ the desired unit quaternion may be expressed as $\mathbf{q}=a_{1}\mathbf{e}_{1}+a_{2}\mathbf{e}_{2}+a_{3}\mathbf{e}_{3}+a_{4}\mathbf{e}_{4}$, with $a_{i}\in\mathbb{R}$.
-The norm $||q||_{\text{quat}}$ can be calculated as $a_{1}^{2}+a_{2}^{2}+a_{3}^{2}+a_{4}^{2}=1$, and now:
+The norm $q||_{\text{quat}}$ can be calculated as $a_{1}^{2}+a_{2}^{2}+a_{3}^{2}+a_{4}^{2}=1$, and now:
 $\mathbf{q}^{T}N\mathbf{q}=\alpha_{1}^{2}\lambda_{1}+\alpha_{2}^{2}\lambda_{2}+\alpha_{3}^{2}\lambda_{3}+\alpha_{4}^{2}\lambda_{4},$ (15)
@@
 One can therefore use a power iteration method or similar algorithm to calculate the principal eigenvalue and eigenvector of $N$ [3].
-II.4 Can we use the Kabsch or Horn algorithms as inspiration for Minkowski vector alignment?
-If we consider minimizing the Frobenius norm of the components, then we follow the Kabsch algorithm in defining the objective function as:
-$f(\Lambda)=\sum_{i}\|v_{B,i}-\Lambda v_{A,i}\|^{2}.$ (22)
-In the Kabsch algorithm, singular value decomposition automatically produces elements of $O(N)$, which allows the simplification step in Eq (7) to Eq (8).
-Further progress using the Kabsch approach would require an efficient numerical Lorentz singular value decomposition [4], which to our knowledge does not exist.
-In the Horn algorithm, the situation becomes more complex (literally).
+∎
+II.2 Can we use the Kabsch or Horn algorithms as inspiration for Minkowski vector alignment?
+For Minkowski vector alignment, we follow the Kabsch algorithm in defining the objective function as:
+$f(\Lambda)=\sum_{i}\|v_{B,i}-\Lambda v_{A,i}\|^{2},$ (16)
+where we again use the Frobenius norm $\|\cdot\|$, thereby aiming to match the numerical values of the 4-vector components between frames $A$ and $B$.
+Proposition 1.
+The Kabsch approach cannot be applied (at present) to minimizing Eq (16).
+Proof.
+In the proof of the Kabsch algorithm, singular value decomposition is the key step which automatically produces elements of $O(N)$, allowing the simplification step in Eq (9) to Eq (10).
+Further practical progress using such an approach would require an efficient numerical Lorentz singular value decomposition [4], which to our knowledge does not yet exist and is beyond the scope of this work.
+∎
+Using the Horn algorithm, the situation becomes more complex (literally).
 Just as unit quaternions represent rotations in $SO(3)$, unit biquaternions represent Lorentz transformations [5].
@@
 Crucially, whereas vectors in $\mathbb{R}^{3}$ were packaged into the imaginary part of the quaternion, 4-vectors must instead be packaged as Minkowski biquaternions, also known as minquats, so that a 4-vector $\mathbf{v}=(t,x,y,z)$ is represented as
-$v=(t,ix,iy,iz).$ (23)
+$v=(t,ix,iy,iz),$ (17)
+where $i$ here is the imaginary unit in $\mathbb{C}$, so that the biquaternion $v$ has purely real (in $\mathbb{C}$) first component and purely imaginary (in $\mathbb{C}$) remaining components.
 Biquaternions have two types of conjugation: complex conjugation in the field, defined by
@@
 The two types of conjugation commute $(\bar{q})^{*}=\overline{(q^{*})}=\overline{q^{*}}$.
-The biquaternion norm is defined as the complex number $||q||_{\text{biquat}}=qq^{*}\in\mathbb{C}$.
+The biquaternion norm is defined as the complex number $q||_{\text{biquat}}=qq^{*}\in\mathbb{C}$.
 The action of the unit quaternion $q$ associated with Lorentz transformation $\Lambda$ on a minquat $v$ is defined by
@@
 The presence of negative contributions to the norm prevents us from writing down the maximum as in Eq (15).
-Schematically, instead of maximizing a quadratic form, we would be maximizing a Hermitian form $\overline{\mathbf{q}}^{T}N\mathbf{q}$, where $N$ is Hermitian and therefore has real eigenvalues $\lambda_{i}$ and orthogonal eigenvectors $e_{i}$, and so that $q$ must be a complex linear combination of the $e_{i}$:
+Proposition 2.
+The Horn approach cannot be applied to minimizing Eq (16).
+Proof.
+Schematically, instead of maximizing a quadratic form as for unit quaternions and $SO(3)$, we would be maximizing a Hermitian form $\overline{\mathbf{q}}^{T}N\mathbf{q}$ for unit biquaternions and $SO(3,1)_{+}$, where $N$ is Hermitian and therefore has real eigenvalues $\lambda_{i}$ and orthogonal eigenvectors $e_{i}$, and so that $q$ must be a complex linear combination of the $e_{i}$:
 $q=\alpha_{1}e_{1}+\alpha_{2}e_{2}+\alpha_{3}e_{3}+\alpha_{4}e_{4},$ (20)
@@
 Unlike in the Horn derivation, we have not linearized the problem at all, and have merely turned one constrained nonlinear optimization problem into another.
+∎
 III Proposed solutions to the Lorentz alignment problem
@@
 III.2 Method 1: Direct nonlinear optimization
-For naive least squares optimization of alignment, we write down the objective function:
+For naive least squares optimization of alignment, we write down the objective function, using the Frobenius norm $\|\cdot\|$ to match numerical components:
 $f(\bm{\zeta},\bm{\theta})=\sum_{i}\|v_{B,i}-\Lambda(\bm{\zeta},\bm{\theta})v_{A,i}\|^{2},$ (31)
-which can be solved either through gradient descent or other method [8]. We opt to use SciPy’s general solvers for this task [9]. We will refer to this method as the direct minimization method.
+which can be minimized either through gradient descent or another method [8], using SciPy’s general solvers for this task [9].
+We will refer to this method as the direct minimization method.
+Note that iterative solvers require a suitable initial value, for which we used the identity matrix.
 III.3 Method 2: Projection onto the Lorentz group after unconstrained solving
@@
 where $A^{+}=(A^{T}A)^{-1}A^{T}$ when $(A^{T}A)^{-1}$ exists [8].
-Consider now the $4\times n$ matrices $X_{ij}=(v_{A,j})_{i}$ and $Y_{ij}=(v_{B,j})_{i}$ (rows index the dimension, and columns index the vectors to be aligned), and the equation
+Consider now the $4\times n$ matrices $X_{ji}=(v_{A,i})_{j}$ and $Y_{ji}=(v_{B,i})_{j}$ (rows index the dimension, and columns index the vectors to be aligned), and the equation
 $\Lambda_{0}X=Y.$ (34)
-The Moore-Penrose inverse gives a least squares estimate for a linear transformation $\Lambda_{0}$ which transforms the vectors $v_{A,j}$ to $v_{B,j}$, but which is not necessarily a Lorentz transform.
+The Moore-Penrose inverse gives a least squares estimate for a linear transformation $\Lambda_{0}$ which transforms the vectors $v_{A,i}$ to $v_{B,i}$, but which is not necessarily a Lorentz transform.
 In other words, we assume that there is some measurement error in $X$ and $Y$ which prevents $\Lambda_{0}$ from satisfying $\Lambda_{0}^{T}\eta\Lambda_{0}=\eta$.
@@
 $l_{0}=\log\Lambda_{0}=\begin{bmatrix}m_{11}&m_{12}&m_{13}&m_{14}\\ m_{21}&m_{22}&m_{23}&m_{24}\\ m_{31}&m_{32}&m_{33}&m_{34}\\ m_{41}&m_{42}&m_{43}&m_{44}\end{bmatrix}$, (35)
-and force $l_{0}$ to be of the form given in Eq (29).
-Specifically, (1) we set the diagonal of $l_{0}$ to be 00, (2) symmetrize the first row and first column, and (3) antisymmetrize the bottom right $3\times 3$ block:
+and consider an objective function of the Frobenius norm
+$f(l)=\|l-l_{0}\|^{2},$ (36)
+$l\in\mathfrak{so}(3,1).$ (37)
+In essence, this objective function forces $l_{0}$ to be of the form given in Eq (23).
+Specifically, (1) we set the diagonal of $l_{0}$ to be 0, (2) symmetrize the first row and first column, and (3) antisymmetrize the bottom right $3\times 3$ block:
 $l=\frac{1}{2}\begin{bmatrix}0&m_{12}+m_{21}&m_{13}+m_{31}&m_{14}+m_{41}\\ m_{12}+m_{21}&0&m_{23}-m_{32}&m_{24}-m_{42}\\ m_{13}+m_{31}&m_{32}-m_{23}&0&m_{34}-m_{43}\\ m_{14}+m_{41}&m_{42}-m_{24}&m_{43}-m_{34}&0\end{bmatrix}$ (38)
-Finally, we take $\Lambda=\exp l$ using Eqs (30)-(35) as the optimal Lorentz transformation. The objective function which this procedure minimizes is:
-$f(l)=\|l-l_{0}\|^{2},$ (43)
-$l\in\mathfrak{so}(3,1)$ (44)
-or
+This solution is seen since minimization of Eq (36) is equivalent to componentwise minimization.
+Since $l\in\mathfrak{so}(3,1)$, the diagonal components of $l$ must be 0, and by symmetry (and anti-symmetry), the off-diagonal elements minimize Eq (36) at the averages of $\bm{\theta}_{i}$ and $\bm{\zeta}_{i}$.
+Finally, we take $\Lambda=\exp l$ using Eqs (24)-(30) as the optimal Lorentz transformation.
+Lemma 1.
+To linear order, minimization in the Lie algebra in Eq (36) and (37) corresponds to minimization of the following function of the Frobenius norm
 $f(\Lambda)=\|\log\left(\Lambda^{-1}\Lambda_{0}\right)\|^{2},$ (39)
 $\Lambda\in SO(3,1).$ (40)
-This method of projection extends easily to all other matrix Lie groups: one simply proposes and minimizes an objective function on the Lie algebra.
+Proof.
+We recall the Baker-Campbell-Hausdorff formula ([10]):
+$\log\left[\exp X\exp Y\right]=X+Y+\frac{1}{2}[X,Y]+\cdots,$ (41)
+where [X,Y]=XYYX[X,Y]=XY-YX. Since $\Lambda=\exp(l)$ and $\Lambda_{0}=\exp(l_{0})$,
+$\log\left(\Lambda^{-1}\Lambda_{0}\right)=\exp(-l)\exp(l_{0})$ (42)
+$\approx-l+l_{0}+\frac{1}{2}[-l,l_{0}]$ (43)
+Theorem 3.
+For sufficiently small measurement noise ($\|\Delta\|\ll\|\Lambda_{\text{GT}}\|$ as below) and for sufficiently small $l,l_{0}]$, minimization using the Lie algebra method corresponds to minimization in the Lie group.
+Proof.
+Let $\Lambda_{\text{GT}}\in SO(3,1)_{+}$ be the ground truth Lorentz transformation, and let there be noise $E$ in the measurement of $Y$:
+$Y=\Lambda_{\text{GT}}X+E.$ (44)
+Using Eq (34),
+$Y=\Lambda_{0}X=\Lambda_{GT}X+E,$ (45)
+Assuming that the Moore-Penrose inverse $X^{+}$ exists,
+$\Delta=\Lambda_{0}-\Lambda_{\text{GT}}=EX^{+}.$ (46)
+If $\|\Delta\|\ll\|\Lambda_{\text{GT}}\|$ for some matrix norm $\|\cdot\|$, then we may write
+$l_{0}=\log(\Lambda_{\text{GT}}+\Delta)=\log\left[\Lambda_{\text{GT}}^{-1}\left(I+\Lambda_{\text{GT}}^{-1}\Delta\right)\right]$ (47)
+$\approx\log\Lambda_{\text{GT}}+\log\left(I+\Lambda_{\text{GT}}^{-1}\Delta\right)$ (48)
+$\approx\log\Lambda_{\text{GT}}+\Lambda_{\text{GT}}^{-1}\Delta-\frac{1}{2}\left(\Lambda_{\text{GT}}^{-1}\Delta\right)^{2}+\cdots$ (49)
+Applying Lemma 1, we have related minimization of Eq (16) in the Lie algebra to minimization in the Lie group, with an error of $O\left(\|\Lambda_{\text{GT}}^{-1}EX^{+}\|+\|[l,l_{0}]\|\right)$. ∎
+To refine the error in Theorem 3, we recall that the Moore-Penrose inverse can be constructed from its singular value decomposition [11], and that the spectral radius of $X^{+}$ therefore is $\frac{1}{\sigma_{\text{min}}(X)}$, where $\sigma_{\text{min}}$ is the lowest magnitude singular value of $X$.
+The factor of $X^{+}$ in the error above therefore contributes a factor of $\frac{1}{\sigma_{\text{min}}(X)}$ to the overall error.
+The $O([l,l_{0}])$ error is nonlinear, but is heuristically low when the projection in Lie algebra space provides a small correction, and can be estimated post hoc.
+We show empirically in the Results that such pathologies were not encountered in what one might consider typical usage of this method.
+If $X$ is rank deficient, one of its singular values must be zero, which may prove a barrier in computing the Moore-Penrose inverse.
+In the case of minimizing Eq (16), this indicates that not enough independent measurements were performed, and there may be infinitely many solutions via the direct minimization method.
+Such an issue can be overcome either by making more measurements or by offsetting the zero-valued singular values by a small real number $\epsilon$.
+The Lie algebra method may fail altogether due to the matrix logarithm failing to exist in Eq (35).
+Heuristically, if the matrix logarithm does not exist, then by continuity of the matrix logarithm and matrix exponential, it must be that the matrix in question was not in the neighborhood of any matrix Lie group, and thus the noise is too large to produce any meaningful results by any method.
+This method of projection extends easily to all other matrix Lie groups: one simply proposes and minimizes an objective function on the Lie algebra, taking care that the appropriate singular values are nonzero, and checking that measurement noise is within tolerable bounds.
+In certain cases the optimal solution in the Lie algebra can be written down explicitly, as in Eq (38).
 Extra consideration is needed if the Lie group is not simply connected and one desires elements other than from the connected component containing the identity.
@@
 The use of the explicit form of the Lorentz boost given by Haber in Eqs (24)-(30) ensured that the resulting matrix indeed had determinant close to $1$.
-However, the Lie algebra method was slightly more effective than direct least squares optimization via the default solver in scipy.optimize.minimize in ensuring that the components of $A\Lambda_{\text{estimated}}^{T}$ were close to $A\Lambda^{T}$, as demonstrated in the Supplemental Materials.
+However, the Lie algebra method was slightly more effective than direct least squares optimization via the default solver in scipy.optimize.minimize in ensuring that the components of $A\Lambda_{\text{est}}^{T}$ were close to $A\Lambda^{T}$, as demonstrated in the attached Code and Data.
 For the Lie algebra method, we tested brute force matrix exponentiation via scipy.linalg.expm compared to Haber’s explicit computation, and found that both computations resulted in $\det\Lambda\approx 1$ to 14 decimal places.
-The timing between the two Lie algebra methods was also extremely similar, with each iteration taking 700-800 $\mu$s, with standard deviation of 20-30 $\mu$s.
-In contrast, the default solver for scipy.optimize.minimize took an average of 24 ms, which was 30$\sim$ 30$\times$ 30 × as slow as the Lie algebra method.
+[…]
 IV.2 Varying the number of vectors given, noise, and averaging over arbitrary Lorentz transformations
-Figure 1: The Frobenius ($L^{2}$) norm of the estimated Lorentz matrix from frame $A$ to frame $B$ as compared to the actual Lorentz matrix.
+In Figs 1 and 2, we compare the absolute Frobenius ($L^{2}$) and max ($L^{\infty}$) norms of the estimated Lorentz transformation matrix ($\Lambda_{\text{est}}$) against the true Lorentz transformation ($\Lambda_{\text{GT}}$).
+We generated $1,000$ random Lorentz transformations using
+$\zeta^{i}\sim\mathcal{N}(0,0.2),$ (52)
+$\theta^{i}\sim\mathcal{N}(0,1).$ (53)
+Figure 1: The absolute Frobenius ($L^{2}$) norm of the estimated Lorentz matrix from frame AA to frame $B$ as compared to the actual Lorentz matrix ($\|\Lambda_{\text{est}}-\Lambda_{\text{GT}}\|_{2}$).
 The rows represent varying the noise $\epsilon$ present in the measurement of the vectors in frame $B$.
 The columns represent varying the number $n$ of vectors upon which the estimation was performed.
-Evidently, the Lie algebra method and the direct minimization method are equivalent in accuracy across a range of testing conditions.
+Evidently, the Lie algebra method and the direct minimization method are equivalent in accuracy across a range of testing conditions, since their absolute error distributions largely overlap.
+Figure 2: The absolute max ($L^{\infty}$) norm of the estimated Lorentz matrix from frame AA to frame $B$ as compared to the actual Lorentz matrix ($\|\Lambda_{\text{est}}-\Lambda_{\text{GT}}\|_{t}ext{max}$).
+The rows and columns represent the same variables as in Fig 1.
 Note that the joint distribution for $\bm{\zeta}$ and $\bm{\theta}$ is a spherically symmetric product of normal distributions.
-We generated $n$ random 4-vectors with Minkowski norm $-1$ by sampling $x,y,z\sim\mathcal{N}(0,0.3)$ and setting $t=\sqrt{1+x^{2}+y^{2}+z^{2}}$, and used these vectors as our $v_{A,i}$.
+We generated $n$ random 4-vectors with Minkowski norm 1-1 by sampling $x,y,z\sim\mathcal{N}(0,0.3)$ and setting $t=\sqrt{1+x^{2}+y^{2}+z^{2}}$, and used these vectors as our $v_{A,i}$.
 We introduced noise in the vector components as measured in frame $B$ by adding random noise $\Delta x,\Delta y,\Delta z\sim\mathcal{N}(0,\epsilon)$ to $x,y,z$, normalizing the time component for $x+\Delta x,y+\Delta y,z+\Delta z$ again, and then boosting by $\Lambda$ to these vectors.
-We observe that the overall distribution of both the maximum element-wise error and the overall $L^{2}$ error is quite similar between the two proposed methods (Figs 1, 2).
-Figure 2: The max ($L^{\infty}$) norm of the estimated Lorentz matrix from frame $A$ to frame $B$ as compared to the actual Lorentz matrix.
-The rows and columns represent the same variables as in Fig 1.
-In Figs 1and 2, we compare the Frobenius and max norms of the estimated Lorentz transformation matrix against the true Lorentz transformation. We generated $1,000$ random Lorentz transformations using
-$\zeta^{i}\sim\mathcal{N}(0,0.2),$ (49)
-$\theta^{i}\sim\mathcal{N}(0,1).$ (50)
+We observe that the overall distribution of both the absolute maximum element-wise error and the overall $L^{2}$ error is quite similar between the two proposed methods (Figs 1, 2).
 V Discussion
@@
 The Horn-like approach is infeasible because of the unbounded nature of the unit biquaternions representation of the Lorentz group, as compared to the bounded real quaternion representation of $SO(3)$; this obstacle prevents the optimization problem from linearizing into a principal eigenvalue/eigenvector problem.
-Instead, we can either perform constrained optimization using gradient descent or related methods (the minimization method) or project the least-squares linear operator via the pseudoinverse onto the Lorentz group via a second least squares minimization in the Lie algebra.
+Instead, we can either perform constrained optimization using gradient descent or related methods (the direct minimization method) or project the least-squares linear operator via the pseudoinverse onto the Lorentz group via a second least squares minimization in the Lie algebra (the Lie algebra method).
 We have shown that at least in Python, the second approach is more performant, and better reproduces the answer in the case that an exact solution exists.
-We did not vary the minimization method used by SciPy’s default solver, though ideally an optimal solution would not be dependent on such as choice.
-The explicit formula relating the boost and rotation vectors to the corresponding Lorentz transformation was necessary in the minimization method, whereas it was optional in the Lie algebra method with little performance difference between it and brute force matrix exponentiation.
-The Lie algebra method presented here technically also solves the orthogonal Procrustes problem by restricting our attention to the bottom right $3\times 3$ matrix.
+We did not vary the direct minimization method used by SciPy’s default solver, though ideally an optimal solution would not be dependent on such as choice.
+The explicit formula relating the boost and rotation vectors to the corresponding Lorentz transformation was necessary in the direct minimization method, whereas it was optional in the Lie algebra method with little performance difference between it and brute force matrix exponentiation.
+The Lie algebra method presented here technically also solves the orthogonal Procrustes problem by restricting our attention to the bottom right $3\times 3$ matrices in the formulae above.
 However, it extends to arbitrary matrix Lie groups, unlike the Kabsch algorithm which only works for $O(N)$, and the Horn algorithm which relies on a quaternion representation with bounded components.
@@
 An application of Lorentz alignment is finding the connection between free-falling reference frames in smooth lattice general relativity [13].
-Future directions on this topic may provide more extensive benchmarking across other matrix Lie groups, and implement the methods described here in a lower level language.
+The direct minimization method requires repeated evaluation of the matrix exponential in Eqs (24)-(30), on the order of tens to hundreds of times, in order to perform gradient descent or some other method of iterative optimization.
+In contrast, the Lie algebra method requires a single evaluation of the matrix logarithm which may be computed iteratively [14], and a single evaluation of the matrix exponential.
+Depending on the convergence speed of the matrix logarithm in comparison to the matrix exponential, the Lie algebra method likely outperforms the direct minimization method for other matrix Lie groups, though further numerical experiments are necessary to verify this hypothesis.
+In both methods, we used an explicit formula to compute the matrix exponential, however this is also not essential since the matrix exponential can be computed through eigendecomposition.
+Again, further numerical experiments would be required to compare the performance of explicit formulae for the matrix exponential versus exponentiation via eigendecomposition for other matrix Lie groups.
+Finally, implementation of these algorithms in other languages such as Julia and C would enable more comprehensive benchmarks.
 Code and data availability
-All code needed to reproduce the data and figures in this work are included in the Supplemental Materials as a Jupyter notebook.
+All code needed to reproduce the data and figures in this work are available at Zenodo (doi: 10.5281/zenodo.15724187).
 Acknowledgments
-The author would like to thank the Penn State College of Medicine’s Medical Scientist Training Program for their support of his career.
+The author would like to thank the Penn State College of Medicine’s Medical Scientist Training Program for its support of his career.
 Conflict of interest statement
 The author reports no potential conflicts of interest.
+Appendix A Quaternions
+Quaternions (often represented by the symbol $\mathbb{H}$) can be thought of as a four-dimensional real vector space with basis $\{\mathbf{1},\mathbf{i},\mathbf{j},\mathbf{k}\}$, with quaternion addition defined by componentwise addition.
+We will represent quaternions as $q=(q_{0},q_{1},q_{2},q_{3})$, where $q_{0},q_{1},q_{2},q_{3}\in\mathbb{R}$.
+We denote the imaginary unit in $\mathbb{C}$ by $h^{2}=-1$.
+The operation of quaternion conjugation is defined by
+$q^{*}=(q_{0},-q_{1},-q_{2},-q_{3}),$ (54)
+Further algebraic properties define the operation of quaternion multiplication:
+$\mathbf{i}\mathbf{j}=\mathbf{k},$ (55)
+$\mathbf{j}\mathbf{k}=\mathbf{i},$ (56)
+$\mathbf{k}\mathbf{i}=\mathbf{j},$ (57)
+$\mathbf{i}^{2}=\mathbf{j}^{2}=\mathbf{k}^{2}=-\mathbf{1},$ (58)
+so that
+$(p_{0},p_{1},p_{2},p_{3})$ $(q_{0},q_{1},q_{2},q_{3})$
+$=($ $p_{0}q_{0}-p_{1}q_{1}-p_{2}q_{2}-p_{3}q_{3},$
+$p_{0}q_{1}+p_{1}q_{0}+p_{2}q_{3}-p_{3}q_{2},$
+$p_{0}q_{2}-p_{1}q_{3}+p_{2}q_{0}+p_{3}q_{1},$
+$p_{0}q_{3}+p_{1}q_{2}-p_{2}q_{1}+p_{3}q_{0}).$ (59)
+Considering $q$ as a column vector, quaternion multiplication $pq$ can also be described as left multiplication of that column vector by a matrix
+$pq=\begin{bmatrix}p_{0}&-p_{1}&-p_{2}&-p_{3}\\ p_{1}&p_{0}&p_{3}&-p_{2}\\ p_{2}&-p_{3}&p_{0}&p_{1}\\ p_{3}&p_{2}&-p_{1}&p_{0}\end{bmatrix}\begin{bmatrix}q_{0}\\ q_{1}\\ q_{2}\\ q_{3}\end{bmatrix}$.
+(60)
+This interpretation of quaternion multiplication as matrix multiplication is used extensively by Horn. Quaternion multiplication is associative but not commutative.
+The quaternion norm is defined as $q||_{\text{quat}}=qq^{*}=q_{0}^{2}+q_{1}^{2}+q_{2}^{2}+q_{3}^{2}$, and unit quaternions by $q||_{\text{quat}}=1$.
+It is unit quaternions which are isomorphic $SU(2)$, which is a double cover of $SO(3)$.
+To represent rotations, a vector $\mathbf{r}\in\mathbb{R}^{3}$ is encoded as a quaternion $r=(0,r_{1},r_{2},r_{3})$.
+We define the operator $U$ which undoes this encoding by $U(r)=\mathbf{r}$.
+Then the rotation of $\mathbf{r}\in\mathbb{R}^{3}$ by a matrix $R\in SO(3)$ can alternatively be performed in quaternion space by:
+$R\mathbf{r}=U(qrq^{*}),$ (61)
+where $q$ is the quaternion representing $R$, the left-hand side $R\mathbf{R}$ is a matrix multiplication, and the right-hand side $qrq^{*}$ is a product of three quaternions, computed using Eq (59).
+The components of $q$ which correspond to a rotation around an axis $\mathbf{u}$ counterclockwise by an angle $\theta$ are:
+$q=\left(\cos\frac{\theta}{2},u_{1}\sin\frac{\theta}{2},u_{2}\sin\frac{\theta}{2},u_{3}\sin\frac{\theta}{2}\right).$ (62)
+Using Eq (62) along with Rodrigues’ rotation formula [6], Eq (61) can be directly verified.